OpenStudy (anonymous):

Write the equation of the line that passes through (2, 4) and (1, –3) in slope-intercept form.

OpenStudy (anonymous):

find slope then use y-b=m (x-a) form, where (a,b)= x,y co-ordinates of your desired poit, example (2,4) or (1,-3)